### v2

# The Forest Behind the Tree: Revealing Hidden Smart Home Communication Patterns

Abstract
The widespread use of Smart Home devices has attracted significant research interest in understanding their behavior within home networks. Unlike general-purpose computers, these devices exhibit relatively simple and predictable network activity patterns. However, previous studies have primarily focused on normal network conditions, overlooking potential hidden patterns that emerge under challenging conditions. Discovering these hidden flows is crucial for assessing device robustness.
This paper addresses this gap by presenting a framework that systematically and automatically reveals these hidden communication patterns. By actively disturbing communication and blocking observed traffic, the framework generates comprehensive profiles structured as behavior trees, uncovering flows that are missed by more shallow methods. This approach was applied to ten real-world devices, identifying 254 unique flows, with over 27% only discovered through this new method. These insights enhance our understanding of device robustness and can be leveraged to improve the accuracy of network security measures.
Index Terms:
IoT, Smart Home, networks, robustness, security, traffic profiling
I Introduction
In the Internet of Things (IoT) paradigm, physical objects are equipped with sensing, computing, and networking capabilities. This allows them to monitor or react to their environment and exchange messages with other objects and with remote servers [1]. A common use case of this paradigm is the Smart Home, composed of household objects, ranging from small power outlets to large appliances. The popularity of Smart Home devices has increased sharply in the past years, with a market size estimated at 101.07 billion USD in 2024 [2]. Their main appeal is to provide home automation, and therefore convenience, to the user.
As IoT devices have very low computing resources, it is a technical challenge to embark them with all the technology necessary for their correct operation. Notoriously, to stay economically competitive, manufacturers tend to expand the user-visible functionalities on the devices’ limited hardware, neglecting transversal aspects such as robustness or security [3, 4]. Ironically, this compromise might impede the initial incentive behind the design of such systems, i.e. user convenience. Indeed, if devices do not provide robustness in the face of network communication instability, the automation system may be unresponsive, penalizing the user. The motivation behind this work is to assess the robustness of Smart Home devices against network instabilities. More precisely, we envision to discover and describe network communication patterns issued by such devices in the case where their default traffic fails.
Due to their goal-oriented nature, Smart Home devices generally exhibit simple, predictable network communication patterns [5]. Consequently, researchers have designed solutions to express their network behavior in a reduced and exhaustive form, a so-called profile. The IETF has even standardized a format for such profiles, the Manufacturer Usage Description (MUD) [6], albeit admittedly relatively limited [7, 8, 9]. Several methods have been presented in the past to automatically create the profile of a device [10, 11, 12]. For this purpose, its network communications are observed over a certain period of time in the wild or under laboratory conditions. Data mining and analysis techniques are then used to extract compact representations of the network traffic and associate them with specific device events. In order to gain a comprehensive view on the communication profile of a device, relatively complex experimental setups are required [13, 14], due to the variety of possible interactions with the device.
We argue that existing approaches to profiling are insufficient for our goal of gaining comprehensive insight into the network behavior of a Smart Home device, as they do not aim at triggering communication patterns that only become visible under certain network conditions, in particular when the default communication mechanism does not succeed. For example, an IoT device might have a list of alternative domain names for its cloud hosted services that it will only contact if the default domain name fails to resolve or the server behind that name does not respond. Ignoring such behaviors leads to a profile, in which some of the device’s communication patterns are missing or incompletely described.
In this paper, we present a framework to comprehensively uncover the communication patterns of a Smart Home device. For each type of interaction with the device, we first observe the traffic flows that occur in an unconstrained network and build a set of flow descriptors from them. Such descriptors contain Internet and transport layer information as well as application layer data, such as domain names. We then iteratively block flows corresponding to certain descriptors and repeat the interaction with the intent of making new flows appear. Deciding which flows to block and the sequence of blocking experiments is not straightforward. Our aim is to discover as many hidden flows as possible but keep the number of experiments under control. To do this efficiently and not end up in an infinite loop where the same patterns appear over and over again, we store the found descriptors in a tree that we traverse using a breadth-first strategy, and prune using an ad hoc heuristic. The result is a set of multi-level, tree-shaped profiles (one for each type of interaction) for the device that describe the device’s default and alternative communication patterns that we then explore to assess the robustness of the devices’ network communications. Our profiles are conceptually similar to the aforementioned ones, e.g. MUD, and can therefore also be used as input configuration for allow-list firewalls. Current state-of-the-art approaches, which do not cover alternative communication paths, might consider the latter as malicious, whereas they are actually part of the device’s intended behavior, only rarer. Profiles obtained from our framework can thus provide more accurate security configurations that do not block the legitimate communications triggered by a robust device.
The contributions of our work can be summarized as follows:
  • We develop and implement a framework, modular and easily extensible, which instruments Smart Home devices and automatically extracts multi-level profiles for them. The core of the framework is a new algorithm that, for a given device interaction, builds an event signature, i.e., the description of flows appearing for that interaction, and then iteratively blocks previously observed flows, until the interaction fails or no new flow is discovered.
  • We apply our framework to a testbed network comprising ten off-the-shelf devices, and generate multi-level profiles for 36 unique events, which sum up to a total of 254 unique network flows discovered, of which 70 were “hidden”, i.e., not part of the default communication patterns of the tested devices, accounting for over 27% of the total number of unique flows.
  • We assess the robustness of the instrumented devices, by computing robustness-related metrics. We conclude that most devices provide at least one backup communication strategy, if a default communication pattern fails.
  • We publish the source code of our signature extraction algorithm and our experimental framework. Those will be made available after the anonymous review process.
The rest of the paper is structured as follows. We describe the problem and the scope of our work in Section II. In Section III, our methodology is explained in detail. Its evaluation is presented in Section IV. Section V presents research works related to ours, and positions them compared to ours. Ultimately, the paper concludes in Section VI.
II Problem statement and scope
Fig. 1 shows a typical Smart Home setup with a smart lamp that can be controlled by a smartphone application (called companion app in the following). The possible network activities in this setup are highly device specific. The device might communicate directly with the companion app in the local home network, or do this via a manufacturer or vendor-operated cloud server (the latter making it also possible to control the device from outside the home network and to receive firmware updates). In fact, many of the devices on the market have both local and remote communication capabilities.
[…]
Figure 1: Typical Smart Home network with a smart lamp
In line with existing work on profiling discussed in Section V, we assume that the network activity is triggered by an interaction with the Smart Home device. The aim of our work is to obtain a profile, i.e., a description of the network communication patterns associated to that device and interaction. We argue that existing approaches on device profiling, that mostly consist in performing the interaction and recording the resulting network activity, are not sufficient to observe all possible communications patterns. Indeed, we expect that some of them only appear when the default (or configured) patterns fail or are disturbed. Our goal is to discover such patterns in an automated way.
The solution that we will present in the next sections requires the collection, inspection and filtering of packets sent to and received from the Smart Home device and the device running the companion app. If the traffic is unencrypted, we leverage its payload. We do not analyze or modify the device firmware, the companion app, or the server side software. We also do not try to decrypt network traffic, therefore we do not consider information in encrypted payload. However, numerous research has shown than relying on unencrypted information and packet or flow metadata, can already provide sufficient information to accurately fingerprint IoT devices [12, 11, 15].
Similarly to related research works, we only consider IP traffic, either wired or wireless. Other protocols exist for Smart Home devices, such as Zigbee [16] and Thread [17] over IEEE 802.15.4 [18], or the proprietary Z-Wave [19]. Such protocols usually rely on a gateway or hub to connect to the local home network. For devices using these protocols, we profile the gateway’s IP traffic. Devices that rely on long-range communication technologies, such as 5G or LoRaWAN [20], are not in the scope of this paper since their traffic cannot be locally filtered or blocked.
Finally, it should be mentioned that devices also communicate for reasons other than being triggered by interactions. Examples include periodic heartbeat messages or messages to discover the local network [21, 14]. A possible approach to profile such activities is to passively monitor the device over a longer period and extract the traffic using periodicity models [14]. Our approach does not pursue such communication activities further, as in this paper we focus on the more complex case of interaction-triggered communication patterns.
III Profiling methodology
In this section, we describe our profiling methodology. We start with an outline of our approach, followed by a detailed presentation of its components.
III-A General workflow
Fig. 2 gives an overview on the workflow of our methodology to obtain a profile for a Smart Home device under a certain type of interaction. An interaction can require the usage of a triggering source, such as a companion app installed on a smartphone. In practice, a Smart Home device may support different interactions (e.g., for a smart lamp: switching on/off, changing its color, etc.) and triggering sources, and the workflow described here must be repeated for each of them. The workflow consists of multiple steps summarized below.
[…]
Figure 2: Overview of the profiling workflow
Step 1 consists in performing the interaction with the device under investigation and recording the network traffic. To this end, we connect the Smart Home device (or, for non-IP devices, their gateway or hub) and the device running the triggering source to a (wireless or wired) LAN and capture the traffic on the Access Point (AP), respectively switch. We call an interaction and the resulting device state an event.
In step 2, the traffic traces obtained are filtered and analyzed, and an event signature is extracted. The latter consists of a set of traffic flow descriptors, called Flow IDs in the following, that describe the relevant features of the bidirectional flows observed in the traces, and therefore associated with the event. A Flow ID contains network and transport layer information, such as host addresses and port numbers, as well as application layer information (when available). The extracted Flow IDs are stored in a tree-shaped data structure that allows tracking which Flow IDs have already been observed.
In step 3, we select a Flow ID from the tree, configure a firewall running on the switch or AP to block packets matching that ID, and repeat steps 1 and 2. The idea is that by deliberately blocking some of the traffic associated with the interaction and repeating the steps 1 to 2, the device and/or app will try alternative communication patterns.
The steps above are repeated until no new patterns are discovered. The final result is the profile of the device for the investigated event. It takes the form of a tree, where each node is a Flow ID linked to the event, and the path from the root towards a given node gives the set of parent Flow IDs which were blocked to let the current Flow ID appear. We define the Flow IDs initially occurring without any traffic blocking required as first-level Flow IDs, whereas the ones appearing as a result of traffic blocking are hidden Flow IDs. The latter are the communication patterns which can be discovered only through our multi-level approach, and therefore missed by state-of-the-art techniques.
III-B Step 1: Traffic capturing
Similarly to PingPong [11] and IoTAthena [15], we start capturing the network traffic, perform the interaction, then stop the traffic capture after a predefined duration $d$. We repeat each interaction and the traffic capturing $m$ times. Our framework automates interactions triggered by a companion app by leveraging the Android Debug Bridge (ADB) tool [22] to generate touch events on the smartphone running the companion app. We also consider power-cycling the Smart home device as an interaction (called boot interaction in the following). To automate the boot interaction, we plug the device in a smart power outlet controlled by our framework.
We collect all incoming and outgoing traffic of the Smart Home device and of the smartphone, since we consider the activities of both of them potentially relevant for the characterization of the event. Between captures, we introduce random waiting intervals to avoid the systematic capturing of periodic background traffic unrelated to the interaction. Control plane packets are filtered out, in particular all ARP, ICMP, DHCP packets, TCP Handshake packets, and TLS Handshake packets, except Client Hello packets containing the Server Name Indication (SNI) extension.
As our approach is based on blocking traffic, it can happen that the interaction does not cause the desired state change in the Smart Home device. For the event signature extraction described in step 2, it is important to filter out such unsuccessful event executions. We do this by verifying the device state, e.g. whether the lamp has turned on, after each capture. To obtain the device state, we leverage the device’s API. We believe this method is more precise and reliable than the approach followed by Mandalari et al. [23] which compares screenshots of the companion app to check whether the displayed status of the Smart Home device has changed. If there is no API, we fall back to screenshots, too, but we compute the Structural Similarity Measure (SSIM) [24] and consider two screenshots identical if the SSIM is over an empirically defined threshold, instead of directly comparing image pixels like Mandalari et al. The result are $m^{+}\leq m$ traffic captures of successful event executions.
III-C Step 2: Event signature extraction
For each of the $m^{+}$ packet traces, the recorded packets are first aggregated to Flow IDs, and then the event signature is extracted. This happens in multiple steps that we now describe.
III-C1 Replacing IP addresses by domain names
During an event, the Smart Home device or the companion app may communicate with cloud services provided by the device vendor. Where possible, we replace non-local source and destination IP addresses by domain names. This makes the signature extraction more robust against changes in the IP addresses caused by cloud migrations or load balancing. To do so, we keep a table matching encountered domain names with their corresponding IP address(es). We first populate the table with entries from the LAN gateway’s DNS cache, and then update it whenever a packet bearing domain data is observed, i.e. a DNS query/response or a TLS Client Hello message with the SNI extension. [^1] We also tried reverse DNS lookups, but since most of the servers contacted during our experiments were hosted at big cloud providers such as Amazon, no meaningful data was obtained, and we abandoned this approach.
III-C2 Aggregating packets to Flow IDs
Packets that share all attributes in the set of properties below are grouped to a bidirectional flow, and the attributes form the Flow ID of that flow.
  • The hostname (domain name or IP address) of the source and destination. The Flow ID’s source is the source of the first packet in the bidirectional flow.
  • The protocol (e.g. TCP or UDP).
  • The source and destination ports. We only consider a port if it belongs to a well known service (e.g. port 80) or if it appears, for the given combination of the other attribute values, in all $m^{+}$ packet traces. This means, for example, that the packets of all TCP connections from a client $A$ to the port 80 of server $B$ are aggregated to a single flow and the random client ports used by $A$ are ignored and not further considered. The same is also done for vendor-specific ports, thanks to the above rule.
  • Application-specific data for known application layer protocols, amongst others: query name and query type for DNS; method and URI for HTTP; message type, code and URI for CoAP. Consequently, for example, a DNS response is grouped with its query (based on the query type and name).
III-C3 Building the event signature
The aggregation of packets to Flow IDs produces a set $f_{i}$ of Flow IDs, with $1\leq i\leq m^{+}$, for each of the $m^{+}$ packet traces. We define the event signature $s$ for the investigated interaction as the set of Flow IDs that appear in all packet traces:
$s:=\bigcap_{i=1}^{m^{+}}f_{i}$
By only keeping the intersection, we filter out Flow IDs belonging to network communications that are not deterministically associated with the event, such as periodic or sporadic messages, as discussed in Section II and subsection III-B.
III-D Step 3: Blocking flows with the event signature tree
The event signature tree is a connected, acyclic, rooted tree where the nodes of the tree are Flow IDs, with the exception of the root node. The tree is used to keep track of the Flow IDs already seen and blocked.
III-D1 Tree creation
At the beginning, when the profiling starts for a specific interaction of the device, the tree only consists of the root node. The firewall running on the switch or AP to which the Smart Home device and the companion app are connected does not block any traffic.
After obtaining the event signature $s$ from step 2, we add the Flow IDs in the signature as a child node to the root node. For the next iteration of steps 1 through 3, we select a node, i.e. a Flow ID, that has not been visited yet from the tree and instruct the firewall to block all packets matching that node and all its ancestors in the tree. The node will be marked as visited and the newly found Flow IDs will be added as children to it, and the procedure is repeated.
The algorithm ends if no unvisited nodes are left. Theoretically, the algorithm could continue to run indefinitely if new Flow IDs are discovered in each iteration. However, in our experiments, the algorithm always terminated after a finite number of iterations because the more Flow IDs are blocked, the more constrained the communication of the device becomes and the fewer successful event traces are captured, until no more new Flow IDs are added to the tree.
III-D2 Tree traversal and pruning
To select the next node, we use a Breadth-First Search (BFS) [25] traversal, i.e., we process all nodes at a given depth before proceeding to the next level. Our objective is to trigger all possible network flows corresponding to a Smart Home device event, i.e., express all possible nodes in the event signature tree. To manage resource usage and execution time, two strategies are used to limit the size of the tree.
Firstly, we already remove event signatures for which less than half of the trace captures were successful, i.e., $m^{+}<m/2$, in step 2. In this way, we filter out event signatures for which there is not enough data to reliably determine the Flow IDs that are with a high probability linked to the event. In practice, we observe a polarization in terms of successful event execution: either all event executions succeed ($m^{+}=m$) or none ($m^{+}=0$). The latter occurs when the firewall rules prevent the event’s success. Setting the threshold at $m/2$ is therefore a conservative choice.
[…]
Secondly, we prune branches of the tree which will likely not provide new information, i.e. no new Flow IDs, by applying an ad hoc pruning heuristic. To determine which heuristic we can apply without losing any information, we ran preliminary experiments with one of our testbed’s devices, the TP-Link HS110 smart plug [26]. We observed that, for any two nodes with the same Flow ID, their children were always identical. We conclude it is likely that no new information will be found by processing a node identical to one which has already been processed, regardless of its position in the tree. Therefore, we adopt the node pruning heuristic, illustrated in Fig. 3: a node is pruned if an equivalent node, i.e., a node with the same Flow ID, has already been processed.
III-D3 How to block packets matching a Flow ID
As explained, a Flow ID comprises traffic features from multiple networking layers, such as hostnames, port numbers, and application layer information. The presence of application layer information in the Flow ID requires a firewall that is able to match and reject traffic based on information found at that layer. A simple firewall, such as the default Linux kernel firewall NFTables [27] (successor of the well-known IPTables) is not sufficient. Instead, we leverage the open-source Smart Home firewall by De Keersmaeker et al. [9]. It is based on NFTables, while enhancing its capabilities to match additional application-layer protocols. However, this firewall can only work with allow rules, while our profiling algorithm is based on blocking rules. Therefore, we modified the firewall, turning it into a deny-list firewall. In our prototype implementation of the profiling algorithm, we translate the list of Flow IDs into firewall blocking rules and let the firewall enforce them for the next batch of experiments. The code of the modified firewall will be made publicly available.
IV Experimental results
In this section, we present our experimental setup, and the results obtained from applying our framework to it.
IV-A Experiment setup
We apply our methodology to a controlled testbed network comprised of real-world devices, mimicking a typical Smart Home network, depicted in Fig. 4. We instrument an array of commercial, off-the-shelf Smart Home devices, including four power plugs, three cameras, and three light bulbs. The profiled interactions, per device category, are the following:
  • For all devices: booting the device;
  • Power plugs: toggling them on/off;
  • Cameras: streaming their video feed;
  • Light bulbs: toggling them on/off, changing their brightness, changing their color.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionLegendWANWired LANWLANZigbee
Figure 4: Experimental Smart Home network.
For all interactions other than booting, we derive up to three usage scenarios, to account for various ways of controlling the device:
  • For all devices, we issue their interactions using their official companion app;
  • For devices that support it, we also control them using the app of the popular home automation platform SmartThings [28];
  • For the TP-Link plug and the Hue light bulb, we also experiment with other apps; respectively, TP-Link Tapo [29] (another app from the same manufacturer, targeted toward Tapo-branded devices) and Hue Essentials [30] (third-party app for Hue devices).
The smartphone running the device-specific apps is a Crosscall CORE-X4 [31], running Android 10.
Device App Interaction Event ID
Power plugs
TP-Link HS110 [26] / boot 1
Kasa Smart toggle 12
TP-Link Tapo toggle 13
SmartThings toggle 17
[…]
SmartThings toggle 14
Tapo P110 [33] / boot 3
TP-Link Tapo toggle 15
SmartThings toggle 11
Woox R5024 (Tuya) [34] / boot 4
Tuya Smart toggle 16
Cameras
Xiaomi MJSXJ02CM [35] / boot 5
Mi Home stream 18
Tapo C200 [36] / boot 6
TP-Link Tapo stream 19
SmartThings stream 21
D-Link DCS-8000LH [37] / boot 7
mydlink stream 20
Light bulbs
[…]
Philips Hue toggle 22
Hue Essentials toggle 23
Hue Essentials brightness 24
Hue Essentials color 25
SmartThings toggle 29
SmartThings brightness 30
SmartThings color 31
Alecto Smart-Bulb10 (Tuya) [39] / boot 9
Tuya Smart toggle 32
Tuya Smart color 33
Tapo L530E [40] / boot 10
TP-Link Tapo toggle 34
TP-Link Tapo brightness 35
TP-Link Tapo color 36
SmartThings toggle 26
SmartThings brightness 27
SmartThings color 28
[…]
The complete device references are listed in Table I, along with the corresponding smartphone apps and interactions instrumented. We only consider IP traffic, whether over Ethernet or Wi-Fi. Therefore, for the two devices communicating using Zigbee, we profile the traffic issued by their hub, i.e. the Hue bridge for the Hue lamp, and the SmartThings hub for the SmartThings Outlet, respectively. For the boot interaction experiments over those devices, we synchronously power-cycle both the device itself and its hub.
Our practical experimental parameters, inspired by related works PingPong [11] and IoTAthena [15], are the following:
  • Number of event iterations: $m=20$;
  • Traffic capture timeout: $d=20$ seconds (90 seconds for event boot);
  • Duration between two event iterations: randomly chosen between 40 (120 for event boot) and 150 seconds.
IV-B Example signature tree
As an example of our framework’s results, we present the event signature tree generated for one device of our corpus, namely the TP-Link HS110 smart plug [26]. This device being a smart power plug, it provides one interaction, i.e. toggling it on and off. For this experiment, we controlled the device using its official Android companion app, i.e. Kasa Smart [41]. The resulting signature tree is displayed in Fig. 5. To avoid unnecessarily cluttering the figure, we only show the first occurrence of each unique flow. In general, every first-level flow occurs as a child of every tree node. In certain cases, a hidden flow will appear at deeper levels, which embodies a backup communication strategy when the default one fails.
[…]
Figure 5: Event signature tree for the TP-Link HS110’s toggle event. Bidirectional flows are represented by a double-headed arrow, whilst a single-headed arrow is used for unidirectional flows. When indicated, the port is associated with its host.
By default, toggling the plug triggers seven different flows:
[…]
    between the phone and the plug’s TCP port $9$;
[…]
    between the phone and the plug’s UDP port $9$;
[…]
    from the phone to the mDNS multicast address (224.0.0.251, UDP port 5353);
[…]
    from the phone to the broadcast address’ (255.255.255.255) UDP port $9$;
[…]
    HTTPS between the phone and the server xx-device-telemetry-gw.iot.i.tplinknbu.com;
[…]
    DNS query/response from the plug to the LAN gateway, for the domain name use1-api.tplinkra.com;
[…]
    HTTPS between the plug and the server use1-api.tplinkra.com.
[…]
IV-C Assessing device event robustness
The experiments in our testbed of devices result in the extraction of 254 unique Flow IDs, of which 70 are hidden ones (i.e. 27.56%). Hidden Flow IDs embody the alternative communication strategies that a device might use if the default one fails, effectively enhancing the device’s robustness to network instability (remember that we only keep Flow IDs associated with successful event executions). As such, we rebrand the count of hidden Flow IDs as the robustness score, and compute it over all the instrumented device events. Out of the 36 instrumented events, 23 (63.9%) have a robustness score of at least one (mean 1.94). This value is encouraging: most types of interactions dispose of at least one backup strategy in the case of a failure. Comprehensive results over all events are shown in the appendix: Fig. 9 shows the count of first-level and hidden Flow IDs; Fig. 10 shows the robustness score.
To gain further insight into the distribution of robustness scores, Fig. 6 shows the same robustness score, grouped along three categories, namely:
  • Device category;
  • Controlling app;
  • Manufacturer.
Figure 6: Robustness score for all device interactions, grouped per device type, app, and manufacturer. Each marker represents the robustness score of one event.
We observe that the different device categories do not provide the same robustness. The most robust devices seem to be the power plugs. A rationale might be that, as the plugs’ operation (switching on or off) is the most simple, it takes less effort to the manufacturers to provide backup strategies.
Regarding the companion app, Tuya seems to be less robust than the others. While most apps provide device control either via the local network or through the cloud, Tuya devices are notoriously reluctant to LAN-based control, preferring cloud-based communication. This decision reduces their robustness by design, as, when the cloud endpoints cannot be reached, the device cannot function properly. SmartThings boasts a good score, both as a controlling app and a manufacturer, helped by a very high robustness score for the SmartThings Outlet’s boot interaction. We can also see that the ”other” app category provides no backup strategy whatsoever, suggesting that a user should prefer using trusted and official apps, even if the devices’ API is theoretically open.
We illustrate the effect of the choice of the app with the Philips Hue lamp device and its toggle interaction, which we have triggered by the official app (Philips Hue), the SmartThings app, and a third-party app (Hue Essentials). The official app and SmartThings show similar results: five and six first-level Flow IDs respectively, and one hidden flow for each. However, when using the Hue Essentials app, only three first-level flows are extracted, and no hidden flow.
Among the manufacturers, the most robust one seems to be Xiaomi, for which the only device in our testbed was a camera. As cameras are usually tied to more critical functions, such as monitoring a home against potential intrusions, Xiaomi followed a respectable rationale here. TP-Link seems quite robust, while Philips, even as a trusted device manufacturer, disappoints in terms of robustness.
IV-D A closer look at DNS data
The Domain Name System (DNS), which allows communicating with hosts by providing a domain name instead of their IP address, is a useful tool to provide application robustness. Here, we steer our analysis towards interactions with the DNS protocol, guided by two questions:
  • If the servers the device tries contacting are unresponsive, will it contact servers with other domain names?
  • If the device’s default DNS resolver is unreachable, will the device try other ones?
We focus on event signatures related to boot interactions, as those proved to provide the most information concerning DNS usage. Fig. 7 shows the data related to both questions: the count of unique domain names contacted, and that of DNS resolvers queried. In both cases, the first-level and the hidden count are shown. The results for all types of interactions is shown in Fig. 11 in the appendix.
[…]
0.4Refer to caption Refer to caption OcFFL
Figure 7: Count of unique domain names / DNS resolvers contacted per device issuing the event boot
IV-D1 Server’s domain names
We observe that, out of the 10 boot events, six consider contacting at least one alternative domain name if the default one fails.
IV-D2 DNS resolvers
Out of the 17 DNS resolvers contacted, only two are hidden resolvers. One of them is simply the LAN gateway (192.168.1.1), meaning the only real new DNS resolver discovered at a level deeper than one is the Chinese public DNS resolver 114.114.114.114, contacted by the Xiaomi camera when the LAN gateway is not responsive. This result shows that, regarding domain name resolving, Smart Home devices still lack robustness. Indeed, every device should try contacting a backup resolver if the primary one fails, as this has been made easy thanks to publicly available resolvers such as Cloudflare’s 1.1.1.1 or Google’s $8.$
IV-E Comparison with related work
In this last part of the evaluation, we compare the profiles generated by our framework with those generated by methods proposed in related work, namely PingPong [11], BehavIoT [14], and MUDgee [42], for the devices present in both their testbed and ours. The former two contributions profile specific individual events, whereas the latter profiles the device as a whole. Table II summarizes the devices and interactions covered by each work.
PingPong BehavIoT MUDgee
[42] [11] [14]
TP-Link HS110 [26] toggle toggle device
Philips Hue lamp [38] toggle device
SmartThings Outlet [32] toggle boot device
D-Link camera [37] stream
TABLE II: Related work device & interaction coverage
Figure 8: Comparison of discovered unique Flow IDs between three related works and ours.
Data from PingPong and MUDgee were taken from their reported results, whereas BehavIoT’s results were reproduced on our side. Fig. 8 compares the count of unique Flow IDs discovered by each work. The three covered related works, by design, do not explore the event signature tree deeper than the first level. Considering event-level profiling, our methodology is, as expected, able to extract more Flow IDs than PingPong and BehavIoT, thanks to a deeper tree inspection, and also a more thorough communication pattern extraction, even at the first level. For device-level profiling, the value representing our work has been computed by merging the event signature trees extracted from all events of the given device. For two of the instrumented devices, our strategy is again able to extract a significantly higher number of Flow IDs than MUDgee. However, the number of Flow IDs is identical for the TP-Link HS110. After manual investigation, it appears the extracted Flow IDs are different: 10 out of 14 MUDgee flows are NTP traffic towards various servers, whereas our solution did not extract any NTP patterns. This traffic might therefore be part of periodic patterns, which are not covered by our strategy. On the other hand, we extract 10 non-NTP traffic patterns which were not detected by MUDgee.
V Related Work
Research on IoT device profiling has been fruitful in the last years. Such works lie at differing layers of abstraction, which can be classified along two orthogonal axes:
  • Object abstraction: which semantic unit is profiled, from individual events, to single devices, to the IoT network as a whole.
  • Behavior abstraction: whether the work intends to profile the low-level network traffic, which can be further dissected into individual packets or aggregate flows, or the higher-level abstract device events retrieved from home automation systems.
In the following, we will summarize related works, and position them along the aforementioned abstraction axes; Table III shows the resulting classification.
Event Device Network
Packets PingPong [11], D-interact [43] IoTa [44] De Keersmaeker [9]
Flows This work MUD [6], Hamza [10], Mandalari [23] De Keersmaeker [9], Girish [4]
Events / BehavIoT [14]
TABLE III: Abstraction levels of device behavior models
Girish et al. [4] conducted the first comprehensive analysis of LAN Smart Home traffic. They exhaustively characterized the protocols used, highlighting the security and privacy vulnerabilities, of intra-network device communication.
Among the numerous proposals for IoT device behavior models, the most widespread is probably the IETF’s MUD standard [6], as it has been standardized and backed up by multiple big players, including Cisco [45]. Subsequently, research leveraging this standard was fueled, including Hamza et al.’s solution to generate MUD profiles from network traces of IoT devices’ traffic [10]. The generated profiles, however, suffer from the shortcomings inherent to the MUD standard, namely, among others, the lack of support for protocols other than IP (v4 & v6) on layer 3, and TCP, UDP and ICMP on layer 4. De Keersmaeker et al. [9] proposed a syntax to express the intended network behavior of devices, inspired by MUD, while overcoming some of its shortcomings, and including support for traffic related to cross-device interactions.
Regarding research efforts to extract event signatures from network traffic, three works follow a workflow similar to ours: PingPong [11], D-interact [43], and IoTa [44]. Nevertheless, all three lie at a finer granularity level than our work, as their signatures are composed of individual packets, and leverage packet-level metadata, such as the packet size. We argue that such features can be subject to network churn and therefore vary from one event execution to the other, making them imprecise to accurately characterize network behavior.
Hu et al. proposed BehavIoT [14], a system which models the behavior of a whole Smart Home network, with the intent of using it to detect when the network produces unintended behavior, potentially due to unwanted communication, or even a security breach. Their model is twofold: on the one hand, they model individual devices events; on the other hand, cross-device interactions. Their complete model is generated based on the network traffic produced by the devices, and takes the form of a probabilistic Finite State Machine (FSM), representing all the possible event transitions in the home network.
All of the above-mentioned works have in common that, unlike our work, they do not actively attempt to discover hidden communication patterns.
VI Conclusion
In this article, we presented a novel methodology and framework to extract multi-level signatures from Smart Home device events, whereas existing works are limited to first-level signatures. Our signatures take the form of trees, one per interaction type with the object, with nodes being the constitutive flows. We show that more than half of the devices exhibit hidden traffic patterns, only occurring when tampering with the network they are connected to. In average, two hidden flows are discovered for each device, across all studied interactions with the device. Our work highlights the robustness capabilities of Smart Home devices, enabling users to take more advised buying decisions. One can envisage that, in the future, devices are labeled with such a robustness score.
As future work, we would like to apply our methodology to scenarios where multiple Smart Home devices are involved in a single interaction. This is conceptually supported by our methodology, but will require an extension of the mechanisms we used for event triggering. In this case, approaches to further automate the latter should also be investigated.